OpenStudy (anonymous):

which answer gives the correct surface area for a triangular prism with bases that are 4cm2 and sides that are 6cm2 ?

jhonyy9 (jhonyy9):

4+4=8 6*3=18 18+8=26 than will be 26 cm squared